OpenStudy (anonymous):

A tree of height y meters has approximately B branches, where B = y − 7. Each branch has approximately n leaves where n = 9(B^(2)) − B. Find the approximate number of leaves of a tree as a function of height.

OpenStudy (aravindg):

\[\huge \color{green}{\text{WELCOME TO OPENSTUDY!}}\]

OpenStudy (anonymous):

Please help?

OpenStudy (tkhunny):

We have \(n = 9B^{2} - B\). This is the number of leaves on each branch. We have \(B = y - 7\). This is the number of branches on a tree. Total Leaves, then, should be \(n*B = (9B^{2} - B)\cdot (y-7)\) Since we are required to find a function of ONLY the height of the tree, we need to get rid of those Bs. This can be done by substitution since we know \(B = y - 7\) \(n*B = (9B^{2} - B)\cdot (y-7) = [9(y-7)^{2} - (y-7)]\cdot (y-7)\) See how that worked? You have some algebra in your future.

OpenStudy (anonymous):

After solving through i got: \[9y ^{3}-187y ^{2}+1337y-3136\] is this my answer or do i need to solve for a value?